OpenStudy (anonymous):

If f(x) = x2 + 3, find f(5). 13 22 28 4

OpenStudy (anonymous):

(5)^2+ 3 = 28

OpenStudy (anonymous):

22

OpenStudy (anonymous):

so who's right??

OpenStudy (anonymous):

I think i got that answer but i dont know who is right

OpenStudy (anonymous):

I'm right...

OpenStudy (anonymous):

Medal me please

OpenStudy (anonymous):

how do u know

OpenStudy (anonymous):

this website is not for medals it is helping others if they have question

OpenStudy (anonymous):

BECAUSE THIS IS 4TH GRADE MATHEMATICS! 5^2 = 25 + 3 = 29

OpenStudy (anonymous):

28